SpaceRef

SpaceRef


SOHO Weekly Meeting Minutes Week 25, June 16 - June 22, 2003

Status Report From: Goddard Space Flight Center
Posted: Monday, June 23, 2003

1. Announcements

----------------
* HGA Anomaly Update: Since resumption of the Z axis movements yesterday,
  we received the first pulse at the expected time, but two pulses expected
  after that have not arrived. 

  Meanwhile, we have also learned that the antenna ground test by
  Saab-Ericsson showed that the center of the beam is not centered on the
  intended zero point of the mechanical coordinates. Thus, the conclusion
  from the 4 June HGA test is now somewhat in doubt - the expected beam
  position for the "motor OK" case can not be determined with sufficient
  precision to support the original conclusion. 

  Nominal operations of the antenna will be continued. If we keep missing
  steps, another offpoint maneuver will be performed to determine the
  antenna movement since 4 June. This should provide a definitive
  discrimination between the "motor OK" and "TM OK" cases. 

  The maneuver is tentatively being scheduled for Wednesday 18 June, with
  Thursday 19 June as a backup. We expect the maneuver to be similar to the
  last one, but somewhat shorter.


2. Operations Constraints
-------------------------
* Commanding times for the week (note: some changes may occur on 
  Wednesday and/or Thursday, if an offpoint maneuver is being performed).

  Monday -    D66 until 17:20 UT then D27 starts at 19:15 UT.
  Tuesday -   D54 until 17:20 UT then D27 starts at 18:15 UT.
  Wednesday - D27 until 21:30 UT then D46.
  Thursday -  D66 until 17:30 UT then D27 until 21:30 UT then D46. 
  Friday -    D27 until 14:30 UT then good stations.
  Saturday -  Commanding is good all day.
  Sunday -    Commanding is good all day.

 

3. Coordinated Observations
---------------------------
M Jun 16 (W25)
T Jun 17
W Jun 18	Offpoint Maneuvers, TBC
T Jun 19
F Jun 20        EIT Shutterless #11 Segment 1 (#6955), 17-19:40 UT
                  EIT/CDS/MDI(TBC)/TRACE/SPIRIT,
                  POC: Jean-Francois Hochedez (j-f.hochedez@oma.be)
S Jun 21
S Jun 22

Other activities for Week 25:
* Week 25       JOP136 Default RHESSI Collaborations (#6850),
                  POC: Stein Haugan
* Week 25       JOP159 CME's in Lyman-alpha (#6870), SWAN/LASCO/EIT
                  POC: Chris St Cyr, Eric Quemerais
* Jun 8-22      Spicules, Mottles and Jets (#6952), 
                  TRACE/La Palma SST/MDI/[CDS], 08-18 UT, POC: Bart DePontieu
* Jun 14-15     UVCS Observations of Zeta Tau
* Jun 14-17     Blinker Study (#6953), MDI(TBC)/La Palma DOT/CDS(TBC), 9-12 UT, 
                  POC: Hardi Peter
* Jun 16-22     UVCS Streamer Study
* Jun 18-23     MDI Continuous Contact

Submodes: Submode 6

Planners for Week 25:
CDS     -- Ben Dobke            MDI     -- Hao Thai/Jeneen Sommers
EIT     -- Matt West            SUMER   -- not observing
LASCO   -- Ops Team             UVCS    -- Leonard Strachan
TRACE   -- Dick Shine


Relevant Notes
--------------

CDS:	Blinker study support, usual weekly studies, EIT shutterless 
	support on Friday.

UVCS:	Streamer study and synoptics.

LASCO:	C2/C3 synoptics, JOP159 Wednesday - Friday. 
 
EIT:	195 CME watch and synoptics, JOP159 Wednesday-Friday,
	shutterless on Friday.

MDI:	La Palma support, blinker study, EIT shutterless support on
	Friday.

TRACE:	La Palma support, EIT shutterless support on Friday.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks